# Gaussian process regression with additive periodic kernels for two-body interaction analysis in coupled phase oscillators

### v1

@@ 4.3 Coupled FitzHugh-Nagumo oscillators
Here is another example which demonstrates the effectiveness of the proposed method. Consider the following coupled FitzHugh-Nagumo oscillators [21, 22]:
@@
We estimated the coupling functions for various numbers of data (Figure 3). Specifically, we varied the number of initial values used to create the data (denoted by $N_{\text{init}}$) and the number of oscillator rounds per initial value used for estimation (denoted by $N_{\text{cycle}}$). We set the sampling rate as 0.2. The phase time series are reconstructed using the same method as described in section 4.2.
From Figure 3, it is evident that when the data sample size is small, our proposed method accurately aligns with the theoretical result in estimating the coupling function, whereas the existing method fails in estimation. This implies that our method is superior to the other methods when a number of the data are small and the data are biased.
Figure 3: Estimated coupling function $\Gamma_{12}$ of the coupled FitzHugh-Nagumo sytem (15). Theory (dotted black curve), results obtained with existing method (solid blue line) and with proposed method (solid red line). (a) $N_{\text{init}}=10,N_{\text{cycle}}=5$. (b) $N_{\text{init}}=30,N_{\text{cycle}}=5$. (c) $N_{\text{init}}=100,N_{\text{cycle}}=5$. (d) $N_{\text{init}}=100,N_{\text{cycle}}=5$ when the part of the data where the phase difference is $0.5\pi$ or more and $\pi$ or less is discarded.
@@ 4.4 Spiking Neural Network oscillators
Finally, we consider a network of multiple coupled Hodgkin-Huxley models [23] and estimate the coupling function for all of them. There are seven neurons, five of which are excitatory and two of which are inhibitory. Details of the Hodgkin-Huxley model is summarized in A.
We employed a kernel function in the form of equation (8) with $N=7$. See the lower graphs of Figure 4 and Figure 5 for the estimation result compared to the theoretical coupling functions obtained by the phase reduction approach. We confirm that the coupling functions between each oscillators are qualitatively consistent with the results from Gaussian process regression and the theoretical results.
Figure 4: (a) A network of spiking neurons coupled with five excitatory and two inhibitory neurons, (b) time series of the action potential of each neuron, (c) theoretically derived coupling functions between each neuron (black dotted lines) and coupling functions estimated from the data using Gaussian process regression (red solid lines: excitatory connection; blue solid lines: inhibitory connection; gray solid lines: no connection). Note that each coupling function $\Gamma_{ij}$ is translated to take 00 at $\Delta\phi_{ij}=0$ for comparison.
Figure 5: Theoretically derived coupling functions between each neuron (black dotted lines) and coupling functions estimated from the data using Gaussian process regression (red solid lines: excitatory connection; blue solid lines: inhibitory connection; gray solid lines: no connection). Note that each coupling function $\Gamma_{ij}$ is translated to take 00 at $\Delta\phi_{ij}=0$ for comparison.
@@ 5 Conclusions and Discussions
In this paper, we focus on the real data that represent rhythmic phenomena, and from them, we consider estimating the coupling function that is modeled as a phase oscillator system. A method for estimating the coupling function using Gaussian process regression is proposed. By designing the kernel function flexibly, the regression can be performed without going through a finite order approximation by Fourier series. We confirmed that this works well in the case of simple phase oscillators, the Van der Pol equations and the FitzHugh-Nagumo equations. We also confirmed that the coupling function can be successfully estimated qualitatively for as many as seven many-body systems of coupled spiking neurons.
Finally, we comment on the proposed method using Gaussian process regression. The main idea was to reduce the data to a regression problem about a vector field of differential equations of phase. However, the possibility arises that the data used in the regression may contain systematic errors because it involves an approximation of the derivative in terms of differences due to finite time. To overcome this problem, a new idea has recently been proposed to calculate the parameter derivatives of time series data obtained from parameterized vector fields using the accompanying equations. We leave the application of the adjoint equation to the estimation of coupling functions as a future work.
Acknowledgements
@@ Appendix A Spiking neurons
We write below for details of the Hodgkin-Huxley model [23].
@@
The current flowing through the synapses, $I_{\mathrm{syn},ij}$, is modeled using the kinetic synapse model [24], where it is represented as
@@
where $T_{ij}$ is the concentration of neurotransmitters, which is set to 1 when a spike is emitted by the presynaptic cell and resets to 0 after 1 millisecond. The constants $\alpha_{ij}$ and $\beta_{ij}$ govern the kinetics of $r_{ij}$, while $E_{ij}$ is the reversal potential and $G_{ij}$ is the synaptic conductance. The values used for excitatory and inhibitory synapses are $(\alpha_{ij},\beta_{ij},E_{ij},G_{ij})=(1.1,0.67,0,0.5)$ and $(9.8,0.2,-75,0.4)$ respectively. Additionally, a weak, independent noise function $\xi_{x,i}$ is added to the membrane voltage $V_{i}$ and channel variables $m_{i}$, $h_{i}$ and $n_{i}$. The noise follows a Gaussian white noise distribution, with $\braket{\xi_{x,i}(t)}=0$ and $\braket{\xi_{x,i}(t)\xi_{y,j}(s)}=\sigma^{2}_{x}\delta_{xy}\delta_{ij}\delta(t-s)$, where $x,y=V,m,h,n$, and $i$ and $j$ are the cell indices. The noise strengths used are $\sigma_{V}=0.5$ and $\sigma_{m}=\sigma_{h}=\sigma_{n}=5\times 10^{-6}$.

### v2

@@ 4.3 Coupled FitzHugh-Nagumo oscillators
Here is another example which demonstrates the effectiveness of the proposed method. Consider the following coupled FitzHugh-Nagumo oscillators FitzHugh1961 ; Nagumo1962 :
@@
We estimated the coupling functions for various numbers of data (Fig. 4). Specifically, we varied the number of initial values used to create the data (denoted by $N_{\text{init}}$) and the number of oscillator rounds per initial value used for estimation (denoted by $N_{\text{cycle}}$). We set the sampling rate as 0.2. The phase time series are reconstructed using the same method as described in section 4.2.
From Fig. 4, it is evident that when the data sample size is small, our proposed method accurately aligns with the theoretical result in estimating the coupling function, whereas the existing method fails in estimation. This implies that our method is superior to the other methods when a number of the data are small and the data are biased.
Figure 4: Estimated coupling function $\Gamma_{12}$ of the coupled FitzHugh-Nagumo system (16). Theory (dotted black curve), results obtained with existing method (solid blue line) and with proposed method (solid red line). (a) $N_{\text{init}}=10,N_{\text{cycle}}=5$. (b) $N_{\text{init}}=30,N_{\text{cycle}}=5$. (c) $N_{\text{init}}=100,N_{\text{cycle}}=5$. (d) $N_{\text{init}}=100,N_{\text{cycle}}=5$ when the part of the data where the phase difference is $0.5\pi$ or more and $\pi$ or less is discarded.
@@ 4.4 Spiking Neural Network oscillators
Finally, we consider a network of multiple coupled Hodgkin-Huxley models h&h and estimate the coupling function for all of them. There are seven neurons, five of which are excitatory and two of which are inhibitory. Details of the Hodgkin-Huxley model is summarized in Appendix A.
We employed a kernel function in the form of equation (9) with $N=7$ and $N=20$. See the lower graphs of Fig. 5 and Fig. 6 for the estimation result compared to the theoretical coupling functions obtained by the phase reduction approach. We confirm that the coupling functions between each oscillators are qualitatively consistent with the results from Gaussian process regression and the theoretical results.
Figure 5: (a) A network of spiking neurons coupled with five excitatory and two inhibitory neurons, (b) time series of the action potential of each neuron, (c) theoretically derived coupling functions between each neuron (black dotted lines) and coupling functions estimated from the data using Gaussian process regression (red solid lines: excitatory connection; blue solid lines: inhibitory connection; gray solid lines: no connection). Note that each coupling function $\Gamma_{ij}$ is translated to take 0 at $\Delta\phi_{ij}=0$ for comparison.
Figure 6: We apply the proposed method to a spiking neural network with 20 neurons (15 excitatory and 5 inhibitory). Theoretically derived coupling functions between each neuron (black dotted lines) and coupling functions estimated from the data using Gaussian process regression (red solid lines: excitatory connection; blue solid lines: inhibitory connection; gray solid lines: no connection). Note that each coupling function $\Gamma_{ij}$ is translated to take 0 at $\Delta\phi_{ij}=0$ for comparison.
@@ 5 Conclusions and Discussions
In this paper, we have proposed a new method for estimating the coupling functions of phase-oscillator models using Gaussian process regression. The kernel functions are designed to capture the characteristics of the coupling functions—additivity and periodicity. We have validated the effectiveness of the proposed method across various oscillator systems, including simple phase oscillators, the Van der Pol equations, and the FitzHugh–Nagumo equations. In particular, we have confirmed that stable estimation is achievable even when the data are biased or limited, with our method outperforming existing approaches. We also have demonstrated that the estimated coupling functions are highly consistent with theoretical ones in many-body system consisting of 20 spiking neurons.
The result in this paper opens new avenues for estimating interactions in many-body systems. Two points are particularly noteworthy. First, our method provides a practical guideline for designing kernel functions, which has long been a challenge in kernel methods. Second, by employing Gaussian process regression, a nonparametric Bayesian inference, our approach is not only robust against data bias but also enables the quantification of uncertainty in the estimation results.
In this study, we focused on rhythmic phenomena, i.e., systems with temporal periodicity. Nevertheless, our approach may also be applicable to systems with spatial periodicity. For example, in a school of swimming fish, as illustrated in the lower left of Fig. 1, interactions that align the orientation (angle) of the fish are expected to occur. In such cases, the interaction function is generally unknown, and researchers have typically assumed simple interactions, such as those in the Vicsek model Vicsek1995 , when constructing models. By contrast, our method enables the estimation of complex interactions directly from data, allowing for the development of models that more accurately reflect reality. In this sense, we believe our findings are highly valuable for applications to many-body systems with unknown interactions, such as active matter.
Acknowledgement
@@ Appendix A Spiking neurons
We write below for details of the Hodgkin-Huxley model h&h .
@@
The current flowing through the synapses, $I_{\mathrm{syn},ij}$, is modeled using the kinetic synapse model Destexhe1994 , where it is represented as
@@
where $T_{ij}$ is the concentration of neurotransmitters, which is set to 1 when a spike is emitted by the presynaptic cell and resets to 0 after 1 millisecond. The constants $\alpha_{ij}$ and $\beta_{ij}$ govern the kinetics of $r_{ij}$, while $E_{ij}$ is the reversal potential and $G_{ij}$ is the synaptic conductance. The values used for excitatory and inhibitory synapses are $(\alpha_{ij},\beta_{ij},E_{ij},G_{ij})=(1.1,0.67,0,0.5)$ and (9.8,0.2,75,0.4)(9.8,0.2,-75,0.4) respectively. Additionally, a weak, independent noise function $\xi_{x,i}$ is added to the membrane voltage $V_{i}$ and channel variables $m_{i}$, $h_{i}$ and $n_{i}$. The noise follows a Gaussian white noise distribution, with $\braket{\xi_{x,i}(t)}=0$ and $\braket{\xi_{x,i}(t)\xi_{y,j}(s)}=\sigma^{2}_{x}\delta_{xy}\delta_{ij}\delta(t-s)$, where $x,y=V,m,h,n$, and $i$ and $j$ are the cell indices. The noise strengths used are $\sigma_{V}=0.5$ and $\sigma_{m}=\sigma_{h}=\sigma_{n}=5\times 10^{-6}$.